The sum of two numbers is 12. Their product is 36. Find the two numbers.
Answer by stanbon(75887)

The sum of two numbers is 12. Their product is 36. Find the two numbers.
-------------
Let one of the numbers be "x".
Then the other number is "12-x".
------------
EQUATION:
x(12-x)=36
12x-x^2 = 36
x^2-12x+36=0
(x-6)(x-6)=0
x=6 (one of the numbers)
12-6=6 (the other number)
================================
Cheers,
Stan H.